Evening!
I am so lost on this problem I don't know how to do this problem please help!

Given m = 3 and the point (-1, 1), which of the following is the point-slope form of the equation?
a.y + 1 = 3(x + 1)
b.y – 1 = 3(x + 1)
c.y - 1 = 3(x - 1)
d.y + 1 = 3(x - 1)

Answer by Alan3354(69443)

I am so lost on this problem I don't know how to do this problem please help!

Given m = 3 and the point (-1, 1), which of the following is the point-slope form of the equation?
---------------
y - y1 = m*(x - x1) where (x1,y1) is the given point and m is the slope
y - 1 = 3*(x +1)